### v2

@@ Abstract
Smoothness is crucial for attaining fast rates in first-order optimization. However, many optimization problems in modern machine learning involve non-smooth objectives. Recent studies relax the smoothness assumption by allowing the Lipschitz constant of the gradient to grow with respect to the gradient norm, which accommodates a broad range of objectives in practice. Despite this progress, existing generalizations of smoothness are restricted to Euclidean geometry with $\ell_{2}$-norm and only have theoretical guarantees for optimization in the Euclidean space. In this paper, we address this limitation by introducing a new $\ell*$-smoothness concept that measures the norm of Hessians in terms of a general norm and its dual, and establish convergence for mirror-descent-type algorithms, matching the rates under the classic smoothness. Notably, we propose a generalized self-bounding property that facilitates bounding the gradients via controlling suboptimality gaps, serving as a principal component for convergence analysis. Beyond deterministic optimization, we establish an anytime convergence for stochastic mirror descent based on a new bounded noise condition that encompasses the widely adopted bounded or affine noise assumptions.
@@ 1 Introduction
This paper considers the convex optimization problem $\min_{\mathbf{x}\in\mathcal{X}}\ f(\mathbf{x})$, where $f$ is a convex differentiable function defined on the domain $\mathcal{X}$. It is well-known that gradient descent converges at the rate of $O(1/\sqrt{T})$ for any Lipschitz continuous $f$ (Nesterov, 2013). If $f$ is smooth, i.e., has a Lipschitz continuous gradient, then a faster rate of $O(1/T)$ can be obtained. Furthermore, an optimal $O(1/T^{2})$ rate is attained by incorporating acceleration schemes (Nesterov, 1983). However, optimization problems arising from modern machine learning (ML) are typically non-smooth (Defazio and Bottou, 2019). Even for standard objectives such as $\ell_{2}$-regression, the global smoothness parameter could still be unbounded (Gorbunov et al., 2025).
Recently, efforts have been made to bridge this gap between theory and practice. Zhang et al. (2020b) propose a generalized $(L_{0},L_{1})$-smooth condition, which assumes $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq L_{0}+L_{1}\left\|\nabla f(\mathbf{x})\right\|_{2}$ for nonnegative constants $L_{0},L_{1}$, on the basis of extensive experimental findings on language and vision models. Under this condition, they analyze the convergence of gradient clipping (Mikolov et al., 2012) and theoretically justify its advantages over gradient descent during neural network training. Later, Li et al. (2023a) propose $\ell$-smoothness which replaces the affine function specified by $(L_{0},L_{1})$ with an arbitrary non-decreasing sub-quadratic function $\ell(\cdot)$, i.e., $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{2})$. The mild condition on $\ell(\cdot)$ ensures compatibility with objectives in modern ML and sheds light on new algorithmic designs.
However, existing studies on generalized smoothness suffer from a critical limitation: the definition is limited to $\ell_{2}$-norm and thus can only relate to optimization algorithms based on gradient descent in the Euclidean space. In particular, none of the existing work has considered mirror descent (Beck and Teboulle, 2003), a powerful optimization algorithm tailored for non-Euclidean problems. Over the years, algorithmic adaptivity to diverse geometries has received tremendous success in traditional ML (Ben-Tal et al., 2001; Zhang et al., 2020c), reinforcement learning (Montgomery and Levine, 2016; Wang et al., 2019; Tomar et al., 2022; Lan, 2023; Alfano et al., 2025), network quantization (Ajanthan et al., 2021), watermarking diffusion models (Liu et al., 2023a), and LLM preference alignment (Munos et al., 2024; Zhang et al., 2025e, d; Wu et al., 2025). Given the indispensable role of mirror descent in these advancements, a natural question arises:
@@
We answer this question affirmatively by devising a new notion of generalized smoothness called $\ell*$-smoothness, which is defined under an arbitrary norm rather than $\ell_{2}$-norm, with full coverage of the $\ell$-smoothness proposed by Li et al. (2023a). To be more specific, we provide a fine-grained characterization of the matrix norm by utilizing a general norm and its dual. Under the $\ell*$-smoothness condition, we establish the convergence of mirror descent, accelerated mirror descent, optimistic mirror descent, and mirror prox algorithms, all matching classic results under $L$-smoothness in deterministic convex optimization. Similar to Li et al. (2023a), the key ingredient in our analysis is to connect the gradients’ dual norm with the suboptimality gap. To this end, we establish a generalized self-bounding property under $\ell*$-smoothness (Lemma 7), which bounds the gradient’s dual norm by the product of the suboptimality gap and the local Lipschitz constant. For different types of mirror descent algorithms, we employ various strategies to demonstrate that the suboptimality gaps along the optimization trajectory can be bounded by the initial one, which may be of independent interest.
Furthermore, we extend our methodology to stochastic convex optimization by proposing a generalized bounded noise condition that significantly broadens the widely used assumption of bounded or affine noise. Inspired by Orabona (2020), we bound suboptimality gaps along the optimization trajectory by converting the high-probability convergence of average-iterate to the last-iterate. Equipped with this powerful tool, we establish an anytime convergence rate of $\widetilde{O}(1/\sqrt{t})$ for stochastic mirror descent. Our main results for both deterministic and stochastic convex optimization is summarized in Table 1. Below, we introduce the notations used in this manuscript.
Notations
[…]
Table 1: Summary of our main results.
Algorithm Convergence Rate* Type**
Mirror Descent (Beck and Teboulle, 2003) $O\left(1/T\right)$ (Theorem 8) a&l
Accelerated Mirror Descent (Lan, 2020) $O\left(1/T^{2}\right)$ (Theorem 9) l
Optimistic Mirror Descent (Chiang et al., 2012) $O\left(1/T\right)$ (Theorem 10) a
Mirror Prox (Nemirovski, 2004) $O\left(1/T\right)$ (Theorem 11) a
Stochastic Mirror Descent (Nemirovski et al., 2009) $\widetilde{O}\left(1/\sqrt{t}\right)$ (Theorem 12) a
    $T$ denotes the total number of iterations, while $t$ represents anytime convergence for any $t\in\mathbb{N}_{+}$.
    This column indicates the types of convergence, either average-iterate (a) or last-iterate (l).
2 Related Work
2.1 $(L_{0},L_{1})$-smoothness
$(L_{0},L_{1})$-smooth: non-convex case
The concept of $(L_{0},L_{1})$-smoothness is firstly proposed by Zhang et al. (2020b) based on empirical observations from LSTMs (Merity et al., 2018) and ResNets (He et al., 2016), which allows the function to have an affine-bounded Hessian norm. Under this new condition, they analyze gradient clipping (Mikolov et al., 2012; Pascanu et al., 2013) for deterministic non-convex optimization and derive a complexity of $O(\epsilon^{-2})$ [^1] In non-convex settings, the goal is to find $\epsilon$-stationary point $\mathbf{x}$ satisfying $\left\|\nabla f(\mathbf{x})\right\|_{2}\leq\epsilon$. In convex settings, we aim at $\epsilon$-suboptimal point $\mathbf{x}$ satisfying $f(\mathbf{x})-f^{*}\leq\epsilon$. The meaning of $\epsilon$ depends of the context., which matches the lower bound (Carmon et al., 2020) up to constant factors. They further extend to stochastic settings and provide an $O(\epsilon^{-4})$ complexity bound under the uniformly bounded noise assumption.
$(L_{0},L_{1})$-smooth: convex case
Koloskova et al. (2023) show that gradient clipping has $O(\epsilon^{-1})$ complexity, matching the classic result in deterministic optimization. Takezawa et al. (2024) establish the same complexity bound for gradient descent with polyak stepsizes (Polyak, 1987). However, besides $(L_{0},L_{1})$-smoothness, these two works further impose an additional $L$-smooth assumption, where $L$ could be significantly larger than $L_{0}$ and $L_{1}$orbunov et al. (2025) address the limitation and recover their results without the extra $L$-smoothness assumption. Moreover, they study a variant of adaptive gradient descent (Malitsky and Mishchenko, 2020) and provide an $O(\epsilon^{-1})$ complexity result, albeit with worse constant terms compared to the original one. For acceleration schemes in convex optimization, they modify the method of similar triangles (MST) (Gasnikov and Nesterov, 2018) and prove an optimal complexity of $O(\epsilon^{-0.5})$.
Other explorations
Ever since Zhang et al. (2020b) proposed $(L_{0},L_{1})$-smoothness condition, this generalized notion of smoothness has been flourishing in minimax optimization (Xian et al., 2024), bilevel optimization (Hao et al., 2024; Gong et al., 2024), multi-objective optimization (Zhang et al., 2025a), sign-based optimization (Crawshaw et al., 2022), distributionally robust optimization (Jin et al., 2021; Zhang et al., 2025b) and variational inequality (Vankov et al., 2024a, b). Numerous attempts have been made to refine existing algorithms under this weaker assumption, including variance reduction (Reisizadeh et al., 2025), clipping/normalized gradient (Zhang et al., 2020a; Qian et al., 2021; Zhao et al., 2021; Hübler et al., 2024; Yang et al., 2024), error feedback (Khirirat et al., 2024b) and trust region methods (Xie et al., 2024a). Notably, Faw et al. (2022, 2023); Wang et al. (2023b); Li et al. (2023b); Wang et al. (2024a); Zhang et al. (2025c) explore the convergence of AdaGrad (Duchi et al., 2011), RMSprop (Hinton et al., 2012), and Adam (Kingma and Ba, 2015) under generalized smoothness. Beyond the optimization community, it has also garnered significant attention in federated learning (Khirirat et al., 2024a; Demidovich et al., 2025), meta-learning (Chayti and Jaggi, 2024), and fine-tuning LLMs (Ma et al., 2025).
2.2 $\alpha$-symmetric Generalized Smoothness
An important generalization of $(L_{0},L_{1})$-smoothness is $\alpha$-symmetric $(L_{0},L_{1})$-smoothness proposed by Chen et al. (2023), which introduces symmetry into the original formulation and also allows the dependency on the gradient norm to be polynomial with degree of $\alpha$. The formal definition is given as follows:
for some $L_{0},L_{1}\in\mathbb{R}_{+}$. For deterministic non-convex optimization, Chen et al. (2023) establish the optimal complexity of $O(\epsilon^{-2})$ for a variant of normalized gradient descent (Nesterov, 1984; Cortés, 2006). They also show that the popular SPIDER algorithm (Fang et al., 2018) achieves the optimal $O(\epsilon^{-3})$ complexity in the stochastic setting. As pointed out by Chen et al. (2023); Vankov et al. (2025), any twice-differentiable function satisfying (1) is also $(L_{0}^{{}^{\prime}},L_{1}^{{}^{\prime}})$-smooth for different constant factors. As an extension, Jiang et al. (2024) study sign-based finite-sum non-convex optimization and obtain an improved complexity over SignSVRG algorithm (Chzhen and Schechtman, 2023).
2.3 $\ell$-smoothness
2.4 Other Generalizations of Classic Smoothness
It is known that a function $f$ is $L$-smooth if $Lh-f$ is convex (Nesterov et al., 2018) for $h(\cdot)=\left\|\cdot\right\|^{2}/2$. Advancements have been made to relax $h$ into an arbitrary Legendre function, leading to the development of a notion called relative smoothness (Birnbaum et al., 2011; Bauschke et al., 2017; Lu et al., 2018; Hanzely et al., 2021; Hanzely and Richtárik, 2021), which substitutes the canonical quadratic upper bound with $f(\mathbf{x})\leq f(\mathbf{y})+\langle\nabla f(\mathbf{y}),\mathbf{x}-\mathbf{y}\rangle+LB_{h}(\mathbf{x},\mathbf{y})$, where $B_{h}$ is the Bregman divergence associated with $h$. Since relative smoothness is now standard in the analysis of mirror descent, we provide a detailed comparison with our $\ell*$-smoothness in Appendix A.
More recently, Mishkin et al. (2024) propose directional smoothness, a measure of local gradient variation that preserves the global smoothness along specific directions. To address the imbalanced Hessian spectrum distribution in practice (Sagun et al., 2016; Pan et al., 2022)Liu et al. (2025); Jiang et al. (2025) propose anisotropic smoothness and obtain improved bounds for adaptive gradient methods.
@@ 3.1 Definitions and Properties
Note that in the Euclidean case, characterizing smoothness either by gradient or Hessian is straightforward due to the nice property $\left\|\cdot\right\|_{2}=\left\|\cdot\right\|_{2,*}$. For the non-Euclidean case where $\left\|\cdot\right\|\neq\left\|\cdot\right\|_{*}$, the previous formulation $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{2})$ in Li et al. (2023a) fails to account for the inherent heterogeneity between the norm and its dual. To clarify, we first rewrite the definition in Li et al. (2023a) into $\sup_{\mathbf{h}\in\mathcal{E}\backslash\left\{\mathbf{0}\right\}}\left\{\left\|\nabla^{2}f(\mathbf{x})\mathbf{h}\right\|_{2}/\left\|\mathbf{h}\right\|_{2}\right\}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{2})$. Then, we underline that the mapping induced by the Hessian matrix (i.e., $\nabla^{2}f(\mathbf{x})\mathbf{h}$) and the gradient (i.e., $\nabla f(\mathbf{x})$) lie in the dual space $\mathcal{E}^{*}$ (Nesterov et al., 2018, Theorem 2.1.6), necessitating the use of the dual norm. To this end, we reformulate the definition into $\sup_{\mathbf{h}\in\mathcal{E}\backslash\left\{\mathbf{0}\right\}}\left\{\left\|\nabla^{2}f(\mathbf{x})\mathbf{h}\right\|_{*}/\left\|\mathbf{h}\right\|\right\}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})$, where the dual norm $\left\|\cdot\right\|_{*}$ is imposed to measure the size of $\nabla^{2}f(\mathbf{x})\mathbf{h}$ and $\nabla f(\mathbf{x})$. The formal definition is given below.
Definition 1
@@
Through careful analysis, we discover that it suffices to take the supremum over the domain $\mathcal{X}$ rather than the whole space $\mathcal{E}$, indicating that Definition 1 enjoys a slightly weaker condition than Li et al. (2023a).
@@
When $\left\|\cdot\right\|=\left\|\cdot\right\|_{2}$efinition 1 simplifies to the $\ell$-smoothness definition in Li et al. (2023a), which recovers the $L$-smooth condition (Nesterov et al., 2018) and recently proposed $(L_{0},L_{1})$-smoothness (Zhang et al., 2020b) with $\ell(\alpha)\equiv L$ and $\ell(\alpha)=L_{0}+L_{1}\alpha$, respectively. In fact, due to the equivalence of norms (Boyd and Vandenberghe, 2004), $\ell$-smoothness can be transformed into $\ell*$-smoothness with a rescaled $\ell(\cdot)$. Whereas, the rescaling brings an additional factor dependent on the problem dimension, ultimately giving rise to dimension-dependent convergence rates.
In general, our formulation encompasses a broader class of functions and also plays a pivotal role in applying mirror-descent-type optimization algorithms. We emphasize that strict twice-differentiability is not required here, as Definition 1 permits (2) to fail on a set of measure zero [^2] One can also follow the similar arguments as in Zhang et al. (2020b) to relax the twice-differentiability condition by considering $\limsup_{\mathbf{h}\to\mathbf{0}}\left\|\nabla f(\mathbf{x}+\mathbf{h})-\nabla f(\mathbf{x})\right\|_{*}/\left\|\mathbf{h}\right\|\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})$.Definition 1 can be seen as a global condition on the Hessian, but it lacks local characterization of the gradients, which proves to be essential for convergence analysis (Zhang et al., 2020b). Therefore, we introduce the following definition to capture the local Lipschitzness of the gradient, which is equivalent to Definition 1 under mild conditions, as illustrated in the sequel.
Definition 1.
($(\ell,r)*$-smoothness) A differentiable function $f:\mathcal{X}\mapsto\mathbb{R}$ belongs to the class $\mathcal{F}_{\ell,r}(\left\|\cdot\right\|)$ w.r.t. a non-decreasing continuous link function $\ell:\mathbb{R}_{+}\mapsto\mathbb{R}_{++}$ and a non-increasing continuous radius function $r:\mathbb{R}_{+}\mapsto\mathbb{R}_{++}$ if (i) $\forall\mathbf{x}\in\mathcal{X},\mathcal{B}(\mathbf{x},r(\left\|\nabla f(\mathbf{x})\right\|_{*}))\subseteq\mathcal{X}$; (ii)
$\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},r(\left\|\nabla f(\mathbf{x})\right\|_{*})),\ \left\|\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2})\right\|_{*}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|.$ (3)
@@
Proposition 2.
(Equivalence between two definitions of generalized smoothness)
  1. (i)
    $\mathcal{F}_{\ell,r}(\left\|\cdot\right\|)\subseteq\mathcal{F}_{\ell}(\left\|\cdot\right\|)$;
  2. (ii)
    under Assumption 1, if $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$, then $f\in\mathcal{F}_{\widetilde{\ell},\widetilde{r}}(\left\|\cdot\right\|)$ where ~$\widetilde{\ell}(\alpha)=\ell(\alpha+G)$ and $\widetilde{r}(\alpha)={G}/{\widetilde{\ell}(\alpha)}$ for any $G\in\mathbb{R}_{++}$.
For simplicity, we restrict our focus to $\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ for the algorithms discussed hereafter. Using 2, we can transform Definition 1 into Definition 1, enabling a quantitative depiction of the local smoothness property around a point $\mathbf{x}\in\mathcal{X}$, as demonstrated in the following lemma.
Lemma 3.
For any $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ satisfying $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G\in\mathbb{R}_{+}$ and given $\mathbf{x}\in\mathcal{X}$, the following properties hold:
  1. (i)
    $\mathcal{B}(\mathbf{x},G/L)\subseteq\mathcal{X}$ where $L:=\ell(2G)$;
  2. (ii)
    $\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},G/L),\ \left\|\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2})\right\|_{*}\leq L\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|$;
  3. (iii)
    $\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},G/L),\ f(\mathbf{x}_{1})\leq f(\mathbf{x}_{2})+\langle\nabla f(\mathbf{x}_{2}),\mathbf{x}_{1}-\mathbf{x}_{2}\rangle+L\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|^{2}/2$.
Lemma 3 provides (i) an optimistic estimate of the local Lipschitz parameter $\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})$, and (ii) functional estimation bounds analogous to $L$-smoothness. This establishes a connection between $\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ and the $L$-smooth function class, demonstrating that the former exhibits similar properties as the latter in a local region. As Lemma 3 requires $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G$, the boundedness of gradients plays a critical role in this transformation which will be discussed further in Section 4.2.

### v3

@@ Abstract
Smoothness is crucial for attaining fast rates in first-order optimization. However, many optimization problems in modern machine learning involve non-smooth objectives. Recent studies relax the smoothness assumption by allowing the Lipschitz constant of the gradient to grow with respect to the gradient norm, which accommodates a broad range of objectives in practice. Despite this progress, existing generalizations of smoothness are restricted to Euclidean geometry with $\ell_{2}$-norm and only have theoretical guarantees for optimization in the Euclidean space. In this paper, we address this limitation by introducing a new $\ell*$-smoothness concept that measures the norm of Hessians in terms of a general norm and its dual, and establish convergence for mirror-descent-type algorithms, matching the rates under the classic smoothness. Notably, we propose a generalized self-bounding property that facilitates bounding the gradients via controlling suboptimality gaps, serving as a principal component for convergence analysis. Beyond deterministic optimization, we establish sharp convergence for stochastic mirror descent, matching state-of-the-art under classic smoothness. Our theory also extends to non-convex and composite optimization, which may shed light on practical usages of mirror descent, including pre-training and post-training of LLMs.
Generalized smoothness, mirror descent, non-Euclidean geometry
Table 1: Summary of our main results.
Algorithm Convergence Rate Type* Convexity
Mirror Descent (Beck and Teboulle, 2003) $O(1/T)$ (Theorem 3.5) a&l
Accelerated Mirror Descent (Lan, 2020) $O(1/T^{2})$ (Theorem 3.7) l
Optimistic Mirror Descent (Chiang et al., 2012) $O(1/T)$ (Equation 11) a
Mirror Prox (Nemirovski, 2004) $O\left(1/T\right)$ (Equation 13) a
Stochastic Mirror Descent (Nemirovski et al., 2009) $\widetilde{O}(1/\sqrt{T})$ (Theorem 4.2) l
Composite Mirror Descent (Duchi et al., 2010) $O(1/T)$ (Theorem 5.1) g
    Types of convergence, either average-iterate (a), last-iterate (l), or gradient-mapping (g).
@@ 1 Introduction
This paper considers the optimization problem $\min_{\mathbf{x}\in\mathcal{X}}\ f(\mathbf{x})$, where $f$ is a convex differentiable function defined on the domain $\mathcal{X}$. It is well-known that gradient descent converges at the rate of $O(1/\sqrt{T})$ for any Lipschitz continuous $f$ (Nesterov, 2013). If $f$ is smooth, i.e., has a Lipschitz continuous gradient, then a faster rate of $O(1/T)$ can be obtained. Furthermore, an optimal $O(1/T^{2})$ rate is attained by incorporating acceleration schemes (Nesterov, 1983). However, optimization problems arising from modern machine learning (ML) are typically non-smooth (Defazio and Bottou, 2019). Even for standard objectives such as $\ell_{2}$-regression, the global smoothness parameter could still be unbounded (Gorbunov et al., 2025).
Recently, efforts have been made to bridge this gap between theory and practice. Zhang et al. (2020b) propose a generalized $(L_{0},L_{1})$-smooth condition, which assumes $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq L_{0}+L_{1}\left\|\nabla f(\mathbf{x})\right\|_{2}$ for nonnegative constants $L_{0},L_{1}$, on the basis of extensive experimental findings on language and vision models. Under this condition, they analyze the convergence of gradient clipping (Mikolov and others, 2012) and theoretically justify its advantages over gradient descent during neural network training. Later, Li et al. (2023a) propose $\ell$-smoothness which replaces the affine function specified by $(L_{0},L_{1})$ with an arbitrary non-decreasing sub-quadratic function $\ell(\cdot)$, i.e., $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{2})$. The mild condition on $\ell(\cdot)$ ensures compatibility with objectives in modern ML and sheds light on new algorithmic designs.
However, existing studies on generalized smoothness suffer from a critical limitation: the definition is limited to $\ell_{2}$-norm and thus can only relate to optimization algorithms based on gradient descent in the Euclidean space. In particular, none of the existing work has considered mirror descent (MD) (Beck and Teboulle, 2003), a powerful optimization algorithm tailored for non-Euclidean problems. Over the years, algorithmic adaptivity to diverse geometries has received tremendous success in traditional ML (Ben-Tal et al., 2001; Zhang et al., 2020c), reinforcement learning (Montgomery and Levine, 2016; Wang et al., 2019; Tomar et al., 2022; Lan, 2023; Alfano et al., 2025), network quantization (Ajanthan et al., 2021), over-parameterization regimes (Sun et al., 2022), watermarking diffusion models (Liu et al., 2023a), pretraining and post-training of LLMs (Xie et al., 2023; Munos et al., 2024; Zhang et al., 2025e, d; Wu et al., 2025). Given the indispensable role of MD in these advancements, a natural question arises:
@@
We answer this question affirmatively by devising a new notion of generalized smoothness called $\ell*$-smoothness, which is defined under an arbitrary norm rather than $\ell_{2}$-norm, with full coverage of the $\ell$-smoothness proposed by Li et al. (2023a). To be more specific, we provide a fine-grained characterization of the matrix norm by utilizing a general norm and its dual. Under the $\ell*$-smoothness condition, we establish the convergence of mirror descent, accelerated mirror descent, optimistic mirror descent, and mirror prox algorithms, all matching classic results under $L$-smoothness in deterministic convex optimization. The key ingredient in our analysis is to bridge the gradients’ dual norm with the suboptimality gap. To this end, we establish a generalized self-bounding property under $\ell*$-smoothness (Lemma 3.4), which bounds the gradient’s dual norm by the product of the suboptimality gap and the local Lipschitz constant. For different types of mirror descent algorithms, we employ various strategies to demonstrate that the suboptimality gaps along the optimization trajectory can be bounded by the initial one, which may be of independent interest.
Furthermore, we extend our methodology to stochastic convex optimization and non-convex composite optimization. For the former, we effectively manage the gradient norms by delving into the last-iterate behavior of stochastic mirror descent (SMD) and provide a simple proof based on the “chain-of-events” (Section 4.2) analysis. We derive a state-of-the-art high probability convergence of $O(\sqrt{\log(T)}/\sqrt{T})$, matching the rate for standard $L$-smooth functions. For the latter, we show that the standard $O(1/T)$ convergence for $L$-smooth functions also holds for $\ell*$-smooth function class. Our main results are summarized in Table 1.
Notations   The optima of the objective is denoted by $\mathbf{x}_{*}:\in\operatorname*{argmin}_{\mathbf{x}\in\mathcal{X}}$ and that $f^{*}:=f(\mathbf{x}_{*})$. Let $\left\|\cdot\right\|$ represent a general norm on a finite-dimensional Banach space $\mathcal{E}$, with its corresponding dual norm defined as $\left\|\mathbf{x}\right\|_{*}:=\sup_{\mathbf{y}\in\mathcal{E}}\left\{\langle\mathbf{x},\mathbf{y}\rangle\mid\left\|\mathbf{y}\right\|\leq 1\right\}$ for any $\mathbf{x}\in\mathcal{E}^{*}$, where $\mathcal{E}^{*}$ is the dual space of $\mathcal{E}$. We denote the ball centered at $\mathbf{x}_{0}$ with radius $r_{0}$ by $\mathcal{B}(\mathbf{x}_{0},r_{0}):=\left\{\mathbf{x}\in\mathcal{E}\mid\left\|\mathbf{x}-\mathbf{x}_{0}\right\|\leq r_{0}\right\}$. The rest notaions can be found in Appendix C.
Related work   We briefly introduce two seminal works on generalized smoothness, with the comprehensive literature review deferred to Appendix AZhang et al. (2020b) first observe in LSTM and ResNet training that the loss functions satisfy an affine bound on the Hessian norm, which they formalize as $(L_{0},L_{1})$-smoothness. Under this assumption, they show that simple gradient clipping finds an $\epsilon$-stationary point in $O(\epsilon^{-2})$ iterations for deterministic non-convex problems and $O(\epsilon^{-4})$ for the stochastic counterpart provided bounded noise. More recently, Li et al. (2023a) propose a broader notion called $\ell$-smoothness, where the Hessian is bounded by a non-decreasing function of the gradient norm. This condition captures a wider range of practical models, and under it, they recover the classic convergence rates: logarithmic in $\epsilon^{-1}$ for strongly convex problems, $O(\epsilon^{-1})$ for convex ones, and $O(\epsilon^{-2})$ for non-convex objectives. They further prove that (modified) Nesterov’s accelerated method achieves $O(\epsilon^{-1/2})$ on convex losses, and that SGD still matches the optimal $O(\epsilon^{-4})$ rate in stochastic non-convex settings (Arjevani et al., 2023).
@@ 2.1 Definitions and Properties
[…]
Definition 2.1.
@@
Through careful analysis, we discover that it suffices to take the supremum over the domain $\mathcal{X}$ rather than the whole space $\mathcal{E}$, indicating that Definition 2.1 enjoys a slightly weaker condition than Li et al. (2023a).
@@
When $\left\|\cdot\right\|=\left\|\cdot\right\|_{2}$Definition 2.1 simplifies to the $\ell$-smoothness definition in Li et al. (2023a), which recovers the $L$-smooth condition (Nesterov and others, 2018) and recently proposed $(L_{0},L_{1})$-smoothness (Zhang et al., 2020b) with $\ell(\alpha)\equiv L$ and $\ell(\alpha)=L_{0}+L_{1}\alpha$, respectively. In fact, due to the equivalence of norms (Boyd and Vandenberghe, 2004), $\ell$-smoothness can be transformed into $\ell*$-smoothness with a rescaled $\ell(\cdot)$. Whereas, the rescaling brings an additional factor dependent on the problem dimension, ultimately giving rise to dimension-dependent convergence rates.
In general, our formulation encompasses a broader class of functions and also plays a pivotal role in applying mirror-descent-type optimization algorithms. We emphasize that strict twice-differentiability is not required here, as Definition 2.1 permits (1) to fail on a set of measure zero [^1] One can also follow the similar arguments as in Zhang et al. (2020b) to relax the twice-differentiability condition by considering $\limsup_{\mathbf{h}\to\mathbf{0}}\left\|\nabla f(\mathbf{x}+\mathbf{h})-\nabla f(\mathbf{x})\right\|_{*}/\left\|\mathbf{h}\right\|\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})$.Definition 2.1 can be seen as a global condition on the Hessian, but it lacks local characterization of the gradients, which proves to be essential for convergence analysis (Zhang et al., 2020b). Therefore, we introduce the following definition to capture the local Lipschitzness of the gradient, which is equivalent to Definition 2.1 under mild conditions, as illustrated in the sequel.
Definition 2.4.
($(\ell,r)*$-smoothness) A differentiable function $f:\mathcal{X}\mapsto\mathbb{R}$ belongs to the class $\mathcal{F}_{\ell,r}(\left\|\cdot\right\|)$ w.r.t. a non-decreasing continuous link function $\ell:\mathbb{R}_{+}\mapsto\mathbb{R}_{++}$ and a non-increasing continuous radius function $r:\mathbb{R}_{+}\mapsto\mathbb{R}_{++}$ if (i) $\forall\mathbf{x}\in\mathcal{X},\mathcal{B}(\mathbf{x},r(\left\|\nabla f(\mathbf{x})\right\|_{*}))\subseteq\mathcal{X}$; (ii) $\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},r(\left\|\nabla f(\mathbf{x})\right\|_{*})),$
$\left\|\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2})\right\|_{*}\leq\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|.$ (2)
@@
Proposition 2.6.
[…]
For simplicity, we restrict our focus to $\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ for the algorithms discussed hereafter. Using Proposition 2.6, we can transform Definition 2.1 into Definition 2.4, enabling a quantitative depiction of the local smoothness property around a point $\mathbf{x}\in\mathcal{X}$, as demonstrated in the following lemma.
Lemma 2.7.
For any $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ satisfying $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G\in\mathbb{R}_{+}$ and given $\mathbf{x}\in\mathcal{X}$, it holds that: (i) $\mathcal{B}(\mathbf{x},G/L)\subseteq\mathcal{X}$ where $L:=\ell(2G)$; (ii) $\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},G/L),\ \left\|\nabla f(\mathbf{x}_{1})-\nabla f(\mathbf{x}_{2})\right\|_{*}\leq L\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|$; (iii) $\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},G/L),\ f(\mathbf{x}_{1})\leq f(\mathbf{x}_{2})+\langle\nabla f(\mathbf{x}_{2}),\mathbf{x}_{1}-\mathbf{x}_{2}\rangle+L\left\|\mathbf{x}_{1}-\mathbf{x}_{2}\right\|^{2}/2$.
Lemma 2.7 provides (i) an optimistic estimate of the local Lipschitz parameter $\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})$, and (ii) functional estimation bounds analogous to $L$-smoothness. This establishes a connection between $\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ and the $L$-smooth function class, demonstrating that the former exhibits similar properties as the latter in a local region. As Lemma 2.7 requires $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G$, the boundedness of gradients plays a critical role in this transformation which will be discussed further in Section 3.2.